Factor a 8x^2 - 6x - 5 = 0 quadratic equation.
1
Expert's answer
2020-08-12T16:18:25-0400

8x2 - 6x - 5 = 0

8x2 + 4x - 10x - 5 = 0


-6x= 4x - 10x We split the middle term -6x to 4x - 10x such that 4*(-10) = -40 which is the same as 8*(-5)=-40.


8x2 + 4x - 10x - 5 = 0

4x (2x + 1) - 5 (2x + 1) = 0

(2x +1) ( 4x - 5)=0


The factors are (2x +1) and ( 4x - 5)